INFRARED SPECTROSCOPY OF BENZENE-HCN IN HELIUM-DROPLETS

\underline{JOCHEN KÜPPER}, KLAAS NAUTA, and ROGER E. MILLER, University of North Carolina, Department of Chemistry, Chapel Hill, NC 27599.

Infrared spectra are reported for several (C_{6}H_{6})_{m}-(HCN)_{n} and (C_{6}D_{6})_{m}-(HCN)_{n} clusters in liquid helium nanodroplets. A large electric field can be used to collapse the rotational structure in these spectra, which aids in the initial searching. In the 1:1 clusters the HCN ν_{CH} stretching vibration at 3253.2 cm^{-1} (C_{6}H_{6}-HCN) and 3252.9 cm^{−1} (C_{6}D_{6}-HCN) are redshifted by approximately 60 cm^{-1} compared to the monomer, indicative of strong intermolecular hydrogen bonds between the CH-group and the aromatic π-clouds of the benzenes. Zero field spectra show considerable rotational structure consistent with a symmetric rotor structure as determined in the gas-phase by microwave spectroscopy.^{a} Moreover several trimers and tetramers of HCN and Benzene have been observed, all of which are further red-shifted than the 1:1 complexes. Pressure-dependency studies allow the size assignments of these clusters. The frequency shifts and rotational contours of the bands allows one to infer the structures of these clusters in liquid helium.

aH.S. Gutowski, E. Arunan, T. Emilson, S.L. Tschopp, and C.E. Dykstra, J. Chem. Phys. 103(1995), 3917.